OpenStudy (anonymous):

Melissa is saving money to buy a game. So far she has saved $30 , which is two-thirds of the total cost of the game. How much does the game cost?

OpenStudy (cwrw238):

30 * 3 / 2

OpenStudy (anonymous):

2/3 of what is 30 2/3x = 30 x = 30/(2/3) x = 30 * 3/2 x = now just multiply like cwrw said